Find the roots of the quadratic equations by the method of completing the square. Express the roots, if they are not rational, in simplest radical form.

Answer by stanbon(75887)

x^2 - 6x - 4 = 0
x^2-6x+9 =4+9
(x-3)^2 = 13
x-3=+-sqrt13
x=3+-sqrt13
x=3+sqrt13 or x=3-sqrt13
Cheers,
Stan H.